OpenStudy (anonymous):

I REALLY NEED SOMEONE TO HELP ME WITH AT LEAST ONE OF THESE PROBLEMS.. PLEASE The solid is similar to a larger solid with the given scale factor, find the surface area and volume of the larger solid. 1) Cylinder: scale factor 1:2, SA=208pi in^2, V=320pi in^3 2)rectangle: scale factor 1:3, SA= 398in^2, V= 495in^3 3) cone: scale factor 3:4, SA=96pi cm^2, V=96pi cm^3

OpenStudy (anonymous):

You would get better help if you post one question per post. You can follow the answers better.

OpenStudy (anonymous):

For the rectangle, you would be finding perimeter and area, are you sure it isn't a box or parallepiped?

OpenStudy (anonymous):

Okay, so Height, Width, Depth V = 495 = HWD S = 398 = 2(DH + HW + DW) Scaling by 3, we get V = 3H*3W*3D = 27HWD=27V = 27*495=1485 S = 2(3D * 3H + 3H*3W+3D*3W) = 2(9DH+9HW+9DW) = 2*9(DH+HW+DW)= 9*2(DH+HW+DW)=9*S=9*398=3582 And of course, dont' forget your units!

OpenStudy (anonymous):

1485 in^3 is the volume, 3582 in^2 is the surface area?

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

yes, and no problem!